Which of the following equations will produce a graph with the intersection of two lines at the positive 2 on the x axis?


A. 3x - 2y =6
2x+y=4
B. -3x-2y=6
-2x+y=4
C. 3x + 2y=6
2x-y= 4
D. 3x - 2y =-6
2x+y = -4


Found 2 solutions by ewatrrr, fcabanski:
Answer by ewatrrr(24785)

 
Hi
two lines intersecting at the positive 2 on the x axis.
That would mean the two lines both have x-intercepts = 2
That is... when y = 0 ⇒ x = 2
A.
3x - 2y =6 || y = 0 ⇒ x = 2
2x + y =4 || y = 0 ⇒ x = 2
C.
3x + 2y=6 || y = 0 ⇒ x = 2
2x - y= 4 || y = 0 ⇒ x = 2

Answer by fcabanski(1391)

One method is to set each equation in the form y=mx+b, set those two equations equal to each other, solve for x, and then plug x back into either equation to solve for y. For this problem you don't have to plug the x value back into one of the equations because the problem doesn't ask for the y value.


A. Equation 1: 3x - 2y =6 (subtract 3x from both sides) -2y =-3x +6 (divide both sides by -2) y = 3/2 x -3


Equation 2: 2x+y=4 (subtract 2x from both sides) y = -2x + 4.


Set the equations = to each other: 3/2 x - 3 = -2x + 4 (add 2x to both sides, add 3 to both sides) 7/2x = 7 (multiply both sides by 2/7) x = 2.


A is an answer because the x value of the intersection point is 2.


B: Equation 1: -3x-2y=6 ---> -2y = 3x + 6 ---> y = -3/2 x - 3


Eq 2: -2x+y=4 ---> y = 2x + 4


Set the equations = and solve for x. -3/2 x - 3 = 2x + 4 ---> 1/2 x = 7 ---> x = 14


B isn't an answer.


C: EQ 1 3x + 2y=6 ---> 2y = -3x + 6 ---> y = -3/2 x + 3


EQ 2 : 2x-y= 4 ---> -y = -2x + 4 ---> y = 2x - 4


Set them equal: -3/2 x + 3 = 2x - 4 ---> -7/2 x = - 7 ---> x = 2


C is an answer.


D: EQ1: 3x - 2y =-6 ---> -2y = -3x -6 ---> y = 3/2 x + 3


EQ2: 2x+y = -4 ---> y = -2x - 4


Set the equations = to each other: -2x - 4 = 3/2 x + 3 ---> -3x = 7 ----> x = -7/3 not an answer.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)